OpenStudy (anonymous):

CAN SOMEONE PLEASE HELP! solve the equation for u t=u+v/2

jimthompson5910 (jim_thompson5910):

Is the equation \[\Large t = \frac{u+v}{2}\] OR is it \[\Large t = u+\frac{v}{2}\]

OpenStudy (anonymous):

just t=u+v divided by 2 and solving for u

jimthompson5910 (jim_thompson5910):

So it's \[\Large t = \frac{u+v}{2}\] right?

OpenStudy (anonymous):